CHWK2-CH1S Problem 15.49 - Enhanced with Feedback Review1 Constants1 Perlodc Table 30 seconds of exposare to 115 dB sound can damage your hearing but a much quieter 94 dB may begin to cause damage aher 1 hour f continuous You are poing to an udoo cancat, and youll be standing neara spett57 Wt
Answer #1

I ( dB) = 10 log ( I / 10^-12)

115 = 10 log ( I/ 10^-12)

I = 0.316 W/m^2

P/ ( 4 pi r^2) = 0.316

57 / (4*3.14* 0.316) = r^2

r = 42.5 m

=====